OpenStudy (anonymous):

can someone please explain this easily:) 4 3/8 + 2 1/2 + 1 3/4

OpenStudy (anonymous):

anyone?

OpenStudy (***[isuru]***):

hi, first turn all the fractions to a common denominator... wt will be the most suitable common denominator here ?

OpenStudy (anonymous):

16

OpenStudy (***[isuru]***):

why didn't u chose 8 ?

OpenStudy (anonymous):

oh yeh

OpenStudy (***[isuru]***):

then let;s change the fractions \[\frac{ 21 }{ 2 }\times \frac{ 4 }{ 4 }=\frac{ 84 }{ 8 }\]

OpenStudy (anonymous):

:) okay

OpenStudy (***[isuru]***):

now do the same for \[\frac{ 13 }{ 4 }\] wt will be the output ?

OpenStudy (***[isuru]***):

\[\frac{ 13 }{4}\times \frac{ 2 }{ 2}=\frac{ 26 }{8 }\] did u get it ?

OpenStudy (anonymous):

yeah thanks

OpenStudy (***[isuru]***):

u r welcome !! hope u can simplify the rest !!!